Calculus, fall 2013

Study Guide for Chap 05 (part 1)

1. Skip section 5.1.

2. Read and understand the meaning of a partition P and its norm ∥P ∥.

3. Read and understand the meaning of a Riemann sum. Practice express a definite integral as the limit of Riemann sum and conversely, practice how to identify a limit of Riemann sum as a definite integral.

4. Study the statement and proof of Fundamental Theorem of Calculus in both Part 1 and Part 2. Try to apply them by making up sum examples to both by yourself.